ARGON-INDUCED PRESSURE BROADENING, SHIFTING AND NARROWING IN THE CN Ã^{2}Π − X̃^{2}Σ^{+} (1-0) BAND

\underline{D. FORTHOMME}, C. P. MCRAVEN, T. J. SEARS^{a}, G. E. HALL, Chemistry Department, Brookhaven National Laboratory, Bldg. 555A, P.O. Box 5000, Upton, NY 11973, USA.

Selected isolated rotational transitions in the 1-0 band of the red Ã^{2}Π − X̃^{2}Σ^{+} system in CN have been recorded with transient frequency modulation spectroscopy as a function of argon pressure up to 0.2 atmospheres at room temperature. Line shapes were fit using Fourier transforms of a parameterized time correlation function, including Doppler and velocitydependent collisional broadening, and collisional shifts. Deviations from Voigt line shapes can be equally well fit by modeling the narrowing with a speed-dependent collision model or with a velocity-changing collisional narrowing model. Pressure broadening coefficients were observed with little rotational state dependence, in the range of 0.070 - 0.075 cm^{−1} atm^{−1}. In contrast, a much stronger rotational state dependence is observed for both pressure-dependent blue shift coefficients and the narrowing parameters. No asymmetry in the pressure broadening was observed; any possible speed-dependence to the frequency shift was too small to be detected in these measurements. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.

aalso : Department of Chemistry, Stony Brook University, Stony Brook, New York 11794, USA